Find the standard of the circle which has a circle at ( 15, -20), radius 9.


1
Expert's answer
2022-01-24T16:25:22-0500

"(15,-20), r=9\\
(x-x_0)^2+(y-y_0)^2=r^2\\
(x-15)^2+(y+20)^2=9^2\\
(x-15)^2+(y+20)^2=81"